y=2x-4
x=1/2y+2
help me to solve this please use elimination method

Answer by stanbon(75887)

y=2x-4
x=(1/2)y+2
help me to solve this please use elimination method
--------------------------
Rearrange:
2x - y = 4
2x - y = 4
-----------------
Subtrct one from the other and you get:
0=0
------------------
That means you have an infinite number of solution points of
the form (x , (1/2)y+2)
------------------------------
Cheers,
Stan H.